# LLM Agents That Act Like Us: Accurate Human Behavior Simulation with Real-World Data

### v4

@@ 4.5 Ablation Study
4.6 Discussion
@@ 5 Conclusion
Our study demonstrates that fine-tuning large language models (LLMs) with real-world human behavioral data and synthesized reasoning traces significantly enhances their ability to generate user actions. These results shed light on the future of LLM agent systems to generate accurate and explainable human-like behaviors in online shopping scenarios and beyond.
6 Ethical Considerations
@@
Appendix A Hyperparameters
We employed Fully Sharded Data Parallel (FSDP, Zhao et al. (2023)) for efficient training. All sequences were padded or truncated to a context length of 40k tokens. We used a per-device batch size of 1, resulting in a global batch size of 64. The learning rate was set to 2e-5 with a cosine scheduler for adaptive learning rate adjustment. Models were trained for 1 epoch. For example, training Mistral-7B-v0.3 with a 40k token context window in our setup requires approximately 130GB of GPU memory per GPU, which is the largest model we can train.
Appendix B Context Definition
The Context, also referred to as the “observation space” of the web agent, encompasses all available information on a webpage, including textual content, metadata, visual elements, and structural data. This context is designed to reflect how a human perceives and interprets a webpage, allowing the agent to process relevant features and perform tasks such as navigation, information retrieval, and interaction with page elements.
Previous research has explored various approaches to defining context. Some methods rely on manually structured information Yao et al. (2022), while others utilize accessibility trees Zhou et al. (2024) or raw HTML representations Gur et al. (2023). However, these approaches have notable limitations: manually structured designs require significant human effort to develop parsing rules Yao et al. (2022), while raw HTML-based approaches often contain extraneous information, such as scripts and raw CSS styles, that are not directly relevant to human users.
To ensure the adaptability to unseen websites, we define and implement a simplified HTML format as our context representation. This format removes non-relevant elements such as scripts, CSS, and purely visual components while preserving essential structural information. Using simplified HTML offers several advantages over custom formats or markdown-based representations: (1) important structural elements, such as lists and tables, remain intact, and (2) LLMs are already familiar with the HTML format, eliminating the need to redefine common elements like ‘button’ and ‘input’ in the prompt.
[…]
{booktabs}
[…]
Llama 3.1 70B 8.19% 12.69%
Mixtral 8x7B 5.41% 13.16%
DeepSeek-R1 11.86% 20.01%
Claude 3.5 Haiku 9.18% 14.77%
Claude 3 Opus 6.78% 15.08%
Claude 3 Sonnet 8.42% 17.40%
Claude 3.5 Sonnet 9.72% 15.91%
Claude 3.5 Sonnet v2 11.69% 18.54%
Claude 3.7 Sonnet 9.34% 12.81%
[…]
w/o reasoning 16.67% 26.92%
Qwen2.5-3B 11.88% 28.52%
w/o reasoning 14.53% 22.88%
Qwen2.5-1.5B 16.06% 27.69%
w/o reasoning 5.03% 5.67%
Mistral-7B-v0.3 15.84% 30.12%
w/o reasoning 14.17% 17.99%
Llama-3.2-3B 15.77% 33.99%
w/o reasoning 9.31% 4.73%
Llama-3.2-1B 7.53% 15.08%
w/o reasoning 11.13% 10.44%
[…]
Table 2: Performance comparison of models in different settings.
[…]
Appendix C Prompts

### v5

@@ 4.5 Ablation Study
4.6 Error Analysis
We analyze model behavior across different error types, comparing two ICL models, Claude, representing general-purpose chat models, and DeepSeek-R1, representing reasoning-augmented models, with a fine-tuned compact model, Qwen2.5–7B. We focus on five distinct error types. Didn’t terminate, didn’t click, and didn’t search indicate cases where the model chose a different action instead of terminating, clicking, or searching, as the human user did. Searched wrong keyword refers to instances where the model performed a search like the user, but used a different (and incorrect) search query. Clicked wrong button captures cases where the model clicked, but on a different button than the one selected by the human user. Illegal actions generated by models are excluded from this analysis. Overall, Claude and DeepSeek-R1 exhibit very similar error profiles, suggesting that the reinforcement learning process used to incorporate reasoning introduces similar biases to those seen in standard chat-based ICL models.
A major shared failure mode of ICL models is their tendency to continue sessions even when human users would have terminated them by closing the browser window. This aligns with earlier findings that ICL agents are more likely to complete purchases, likely because current LLMs are optimized to fulfill task goals rather than follow subtle social cues or termination heuristics. In contrast, Qwen2.5–7B demonstrates more accurate alignment with human termination behavior.
We also find that Qwen2.5–7B better captures iterative search behavior commonly seen in human users, such as retrying search with a different search keyword after an unsatisfactory result or correcting typos in queries. In comparison, ICL models often stick to the original search result and conduct another type of action when a real human user would have made another search. These findings suggest that fine-tuning not only reduces the overall error rate but also enhances the model’s ability to capture fine-grained user behavior patterns, such as correcting typos, retrying failed searches, and deciding when to terminate a session—behaviors commonly observed in real users but often overlooked by ICL models.
5 Discussion and Future Works
5.1 Action Misalign Between Human and LLM Agents
@@
Building on this observation, we find that current LLM agents are often behaviorally misaligned with real users, particularly in the shopping domain. For example, as shown in Figure 2 and Figure 3, out-of-the-box LLMs tend to make more purchases than human users. We hypothesize that this stems from the training objectives of many LLM agents, which are typically optimized for task completion (e.g., sending emails) and evaluated based on completion rates and task efficiency. To enable accurate human behavior simulation, it is essential to close this behavioral gap and ensure models capture not just goal-oriented actions, but also the nuanced, often inefficient behaviors of real users.
The ability to simulate nuanced user behavior, including exploratory actions and natural errors such as typos, opens up exciting new research directions. For example, in automatic UI/UX evaluation Lu et al. (2025b), simulating realistic user behavior is crucial: to evaluate a typo-correction feature, an LLM agent must be capable of making typos in a human-like way. Such behavioral fidelity is difficult to achieve without training models on fine-grained, human-aligned data.
5.2 Reasoning in Next Action Generation
@@ 6 Conclusion
In this work, we present the first quantitative, process-centric evaluation of state-of-the-art large language models (LLMs) for simulating human behavior in an online shopping task. Our study demonstrates that fine-tuning LLMs with real-world human behavioral data and synthesized reasoning traces significantly enhances their ability to generate user actions. Compared to in-context learning and pretrained baselines, our fine-tuned models achieved substantial improvements, including a 13.98% absolute gain in F1 score for final outcome prediction and a 5.4% increase in action generation accuracy.
These results underscore the critical role of explicit reasoning traces in aligning model predictions with human behavior. By enriching behavioral datasets with structured reasoning, we move closer to accurate and interpretable simulations of human behavior. Our findings highlight a promising direction for the development of LLM agent systems capable of producing realistic and explainable human-like behaviors in online shopping and other interactive domains.
Limitations
@@
Appendix A Prompts
@@
Appendix B Example Context
[…]
[breaklines,breakafter=_<>]htmlexample_data.html